OpenStudy (anonymous):

Find the domain of f \[f(x)=(2x+7)^{1/2}\]

OpenStudy (anonymous):

I know how to solve the inequality to get \[x \ge -\frac{7}{2}\] Never mind I got it... \[[-\frac{5}{7}, \infty)\]